Re: Shift report 20000810 00:00-08:00

From: Peter H. L. Christiansen (pchristi@nbi.dk)
Date: Thu Aug 10 2000 - 11:51:20 EDT

  • Next message: http account: "Shift report 20000810 08:00-16:00"

    > > 2.  There were tracks in all tpc.  2 Tiles for calibration 
    > > were read out.
    
    I found the tracks myself and they look real pretty.
    
    I looked at:
    run002181seq001.dat
    where event 9655 and 10295 both have 2 tracks ( 1 hard and 1 weak( read :
    turn down adc cut) ). In both cases there are 2+ tracks in T1.
    
    Looks great!
       Peter 
    



    This archive was generated by hypermail 2b29 : Thu Aug 10 2000 - 11:54:47 EDT